OpenStudy (anonymous):

I REEEEEAAAAAAALY need help :c Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (nuke_ur_friend):

do u like mlp

OpenStudy (anonymous):

Yes I do c:

OpenStudy (nuke_ur_friend):

imma a brony

OpenStudy (arthur_ser):

y=mx + b 6=(-3)(-1) + b b=6-3 b=3 you know m=-3 and now you know b

OpenStudy (nuke_ur_friend):

so u are awesome my friend

OpenStudy (anonymous):

Thank you @arthur_ser heheh, thanks, You're cool too @Nuke_ur_friend